Grade 12th passMechanics

a ball is thrown vertically upwards from the ground if T1 and T2 are the respective time taken in going up and coming down and the air resistance is not ignored then
a] T1>T2
b] T1
c] T1=T2

ishita bist
10 Years agoGrade 12th pass

1 Answer

PASUPULETI   GURU MAHESH
10 Years ago
up ward ball      initial  velocity= u
                         final velocity   =v =0
  
   we know that v= u -gt   (since ball going up ward)
          then t= u/g
 
down ward ball   initial  velocity= u =0
                         final velocity   =v
 
we know that v= u+gt
           then t=v/g
 
by chose one example then the  t1=t2